Calculate no of electrons in 10 grams of water &1 mole of oxygen

Grade:11

1 Answers

Umakant biswal
6 years ago
dear ubi 
no of moles of wtaer = given mass / molar mass = 100 / 18 = 50 / 9 moles 
total no of atoms in 100g of water = 3.455 * 10 ^ 24 = avg no * no of moles 
no fo electrons in 1 atom of h20 = 2 * 1 +8 = 10 
total no of electrons = 3.455 * 10 ^ 25 
charge in 1 electron = 1.6 * 10 ^ -19 
charge carried by 3.455 * 10 ^ 15 electrons = 5.528 * 10 ^ 6 
HOPE IT CLEARS NOW 
REGARDS